# Memory-Enhanced Neural Solvers for Efficient Adaptation in Combinatorial Optimization

### v2

@@ 4 Experiments
We evaluate our method across widely recognized CO problems Travelling Salesman (TSP) and Capacitated Vehicle Routing (CVRP). These problems serve as standard benchmarks for evaluating RL-based CO methods (Deudon et al., 2018; Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023; Hottung et al., 2022; Choo et al., 2022; Chalumeau et al., 2023b). In Section 4.1, we validate our method by comparing it to leading approach for single-policy adaptation using online collected data, EAS (Hottung et al., 2022). We present their comparative performance in line with established standard benchmarking from literature, and elucidate the mechanisms employed to adapt the action distribution of the base policy during inference. We then demonstrate how MEMENTO can be combined with SOTA neural solver COMPASS with no additional retraining in Section 4.2. Finally, we scale those methods to larger instances in Section 4.3, outperforming heatmap-based methods.
@@
Figure 3: Results of MEMENTO against baselines on a standard benchmark comprising $8$ datasets with $4$ distinct instance sizes on TSP (a) and CVRP (b). Methods are evaluated on instances from training distribution ($100$) as well as on larger instance sizes to test generalization. The right panel aggregates those results, reported in % improvement relative to a greedy rollout of the given base policy (POMO). MEMENTO is consistently outperforming policy-gradient based method EAS and tree-search based method SGBS, with significant improvement over the base policy. *SGBS results are reported from Choo et al. (2022), inducing biases in the comparison, discussed in Section A.1.
@@
We provide access to the code [^1] Code, checkpoints, and evaluation sets at https://github.com/instadeepai/memento utilized for training our method and executing all baseline models. Additionally, we release our checkpoints for all problem types and scales, accompanied by the necessary datasets to replicate our findings. We implement our method and experiments in JAX (Bradbury et al., 2018), in a codebase that is compatible with previous JAX implementation released in the community. The two problems are also JAX implementations from Jumanji (Bonnet et al., 2023), enabling us to fully leverage hardware accelerators such as TPUs. Our implementations are optimized for TPU v3-8, aligning with the hardware utilized in our experiments.
@@ 4.1 Benchmarking MEMENTO against policy gradient fine-tuning
The most direct way to leverage online data for policy update is RL/policy-gradients. By contrast, MEMENTO learns an update based on the collected data. We therefore want to see how those two approaches compare. Since EAS (Hottung et al., 2022) is SOTA method using policy gradient, we benchmark MEMENTO against it on a standard set of instances used in the literature (Kool et al., 2019; Kwon et al., 2020; Hottung et al., 2022). Specifically, for TSP and CVRP, those datasets comprise 10 0001000010\,00010 000 instances drawn from the training distribution. These instances feature the positions of $100$ cities/customers uniformly sampled within the unit square. The benchmark also includes three datasets of distributions not encountered during training, each comprising of $1000$ problem instances with larger sizes: $125$, $150$, and $200$, generated from a uniform distribution across the unit square. We employ the exact same datasets as those utilized in the literature for consistency and comparability.
@@
For routing problems like TSP and CVRP, POMO (Kwon et al., 2020) is the base single-agent, one-shot architecture that underpins most RL construction solvers. In this set of experiments, MEMENTO and EAS both use POMO as a base policy, and adapt it to get the best possible performance within a given budget. We train MEMENTO until it converges, on the same instance distribution as that used for the initial checkpoint. When assessing active-search performance, each method operates within a fixed budget of $1600$ attempts, a methodology akin to Hottung et al. (2022). In this setup, each attempt comprises of one trajectory per possible starting point. This standardized approach facilitates direct comparison to POMO and EAS, which also utilize rollouts from all starting points at each step. Following recent works, we do not use augmentations with symmetries which have proved to not be critical and greatly increase the cost (Chalumeau et al., 2023b).
Figure 4: Heatmaps of the update rules induced by REINFORCE gradient (left), compared to those learned by MEMENTO (right). Akin to REINFORCE, MEMENTO encourages actions with high returns, particularly when they have low probability. Interestingly, MEMENTO learns an asymmetric rule: requiring the normalised return to be strictly positive to reinforce an action, but encouraging it even more when that condition is reached (higher amplitude for positive updates than negative ones).
@@ Results
The average performance of each method across all problem settings are presented in Fig. 3. The observations we draw are three-fold. First, MEMENTO outperforms its base model (POMO) on the entire benchmark by a significant margin: showing that its adaptive search is superior to stochastic sampling. Second, MEMENTO outperforms EAS on all $8$ tasks (spanning both in- and out-of-distribution) for both environments, highlighting the efficacy of learned policy updates when compared to vanilla policy gradients. Finally, we note that this improvement is significant; for example, on TSP100, MEMENTO is doing $60$% better than sampling, while EAS only $6$%.
@@
To better understand how MEMENTO uses its memory to adapt the action distribution of its base policy, we analyse the logit update that an action gets with respect to its associated data, akin to similar analysis conducted in meta-learning (Lu et al., 2022; Jackson et al., 2024), and compare it to the REINFORCE policy gradient. On Fig. 4, we plot the heatmap of the logit update with respect to the log probability (logp) and return associated with an action. We observe that the main rules learned by MEMENTO are similar to REINFORCE: an action with low logp and high return gets a positive update while an action with high logp and low return gets discouraged. It is nonetheless interesting to see discrepancies. In particular, MEMENTO’s rules are not symmetric with respect to $x=0$, it only encourages action that are above the mean return, but gives higher amplitude to high-return actions than it discourages low-return ones, pushing for performance while preserving exploration. A similar dissymetry would be expected with EAS since it combines REINFORCE with a term that increases the likelihood to generate the best solution found.
Additionally, MEMENTO uses more inputs than REINFORCE: current budget available, trajectory logp, age of the data, and previous MEMENTO logit are also used to derive the new action logit. Appendix G provides an ablation study of those additional inputs, validating their importance.
4.2 Zero-shot combination of MEMENTO with unseen solvers
MEMENTO was designed to be agnostic to the base model, enabling combination with existing and future solvers. In this section, we demonstrate zero-shot combination (no re-training) with state-of-the-art solver COMPASS. COMPASS provides fast adaptation since it uses a collection of diverse pre-trained policies that can be searched to pick the most relevant policy for unseen instances.
Figure 5: Evolution of performance over budget when combining MEMENTO to COMPASS on CVRP200, no re-training needed.
Although providing fast adaptation, it can hardly improve further once the appropriate pre-trained policy has been found. In particular, online collected data cannot be used to update that particular policy’s action distribution. We use MEMENTO’s memory module, trained with POMO as a base policy, and apply it on COMPASS with no additional re-training, and show empirically that we can switch on MEMENTO’s adaptation mechanism during the search and reach a new state-of-the-art on $11$ out of $12$ tasks.
@@ 4.2 Zero-shot combination of MEMENTO with unseen solvers
We combine our checkpoint of MEMENTO with a released model of COMPASS (Chalumeau et al., 2023b). For the first half of the budget, we do not use MEMENTO, to let COMPASS find the appropriate pre-trained policy. Once the search starts narrowing down, we activate MEMENTO to adapt the pre-trained policy, with no need to turn off COMPASS’ search. Fig. 5 shows the typical trends observed at inference. We see the typical search reported in Chalumeau et al. (2023b). At $50$% of budget consumption, MEMENTO is activated and one can observe a clear and significant drop in averaged tour length of the latest sampled solutions. The standard deviation also illustrates how the search is suddenly narrowed down.
Table 1: Zero-shot combination of MEMENTO and COMPASS on CVRP. Rules learned by MEMENTO transfer to population-based method COMPASS and achieve SOTA on standard benchmark.
@@
Results
We evaluate the zero-shot combination of MEMENTO and COMPASS on the standard benchmark reported in Section 4.1. Results for CVRP are presented in Table 1, showing tour length and optimality gap in-distribution (CVRP100) and out-of-distribution. Although requiring no re-training, MEMENTO’s rules combine efficiently with COMPASS and achieves new SOTA. Results on TSP are reported in Appendix A, and results on larger instances (TSP & CVRP) in Section 4.3.
@@ 4.3 Scaling RL construction methods to larger instances
Scaling neural solvers to larger instances is a crucial challenge to be tackled in order to bridge the gap to industrial applications. Auto-regressive construction-based solvers are promising approaches, requiring limited expert knowledge whilst providing strong performance. But are still hardly ever trained on larger scales with RL. As a result, their performance reported in the literature is usually quite unfair, being evaluated in large scales instances (>$500$) although having been trained on $100$ nodes. In this section, we explain how we have trained construction solvers POMO and COMPASS on instances of size $500$ with RL (for both TSP and CVRP, checkpoints to be open-sourced); and then use them as new base models to train and validate properties of MEMENTO at that scale.
Figure 6: Performance improvement of MEMENTO over EAS on instances of size $500$ on TSP (left) and CVRP (right) for increasing number of sequential attempts and sizes of attempt batch. Green areas correspond to settings for which MEMENTO’s online adaptation is more efficient than EAS. MEMENTO always outperforms EAS on TSP, and for most settings on CVRP. The difference is particularly significant for low budgets, showing the data-efficiency and robustness of MEMENTO.
@@
Training POMO on instances of size $500$ involves three main steps. First, inspired by curriculum-based methods, we start from a checkpoint pre-trained on TSP100. Second, we use Efficient Attention (Rabe & Staats, 2022) to reduce the memory cost of multi-head attention, enabling to rollout problems in parallel despite the $O(n^{2})$ memory requirement. Third, we use gradient accumulation to keep good estimates despite the constrained smaller instance batch size. These three combined tricks enable to train POMO till convergence within $4$ days, and consequently to train COMPASS and MEMENTO. We also build the zero-shot combined MEMENTO (COMPASS) checkpoint. We release all checkpoints (TSP & CVRP) publicly to ease future research at this scale.
@@
On those larger instances, the cost of constructing a solution is significantly increased (for POMO’s architecture: squared in computation and memory, and linear in sequential operations). It is also harder to get numerous parallel shots, which most methods depend on. It hence becomes crucial to develop data-efficient methods which stay robust to low budget regimes; and important to know which method to use for each budget constraint.
@@
We compare MEMENTO and EAS on a set of 128 unseen instances (of size $500$), for a range of budget expressed in number of sequential attempts and number of parallel attempts attempts. Figure 6 reports the percentage of improvement (in absolute cost) brought by MEMENTO over EAS. From this figure, we can observe three main tendencies. First, the data-efficiency of MEMENTO is illustrated by its consistent superiority on low budget (lower left of plots), reaching significant improvement compared to EAS. Then, we can see that, on CVRP, the gap increases with the sequential budget for low parallelism. It demonstrates that MEMENTO’s update stays robust, whereas the gradient estimates of EAS get deteriorated and cannot bring further improvement despite additional sequential budget. To finish with, we can see that MEMENTO fully dominates the heatmap on TSP, even for higher budget. Nevertheless, when budget for CVRP becomes bigger, EAS is able to outperform MEMENTO, given that batches are large enough ($>80$).
(a) TSP (n=500)
Method Obj. Gap Time -
Concorde 16.55 $0.000\%$ 38M -
LKH3 16.55 $0.003\%$ 46M -
Dimes 17.80 $7.55\%$ 2H -
Difusco 17.23 $4.08\%$ 11M -
DeepACO 18.74 $13.23\%$
MOCO 16.84 $1.72\%$
Pointerformer 17.14 $3.56\%$ 1M -
Low Budget High Budget
Method Obj. Gap Time Obj. Gap Time
@@
12M
@@
36M
(b) CVRP (n=500)
Method Obj. Gap Time -
LKH3 37.229 $0.00\%$ 6H -
Low Budget High Budget
Method Obj. Gap Time Obj. Gap Time
@@
15M
@@
47M
Figure 7: Tables on the left report results of MEMENTO against the baseline algorithms on instances of size $500$ of TSP (a) and CVRP (b) for two budget regimes (low and high). We observe that POMO-based methods are outperforming concurrent methods. MEMENTO achieves single-agent SOTA (resp. overall SOTA) on $3$ out of $4$ settings when combined with POMO (resp. COMPASS). The plots on the right show the time complexity of MEMENTO and EAS for increasing values of decoder size, and instance size. It demonstrates that MEMENTO scales better in time than EAS.
Results
In Fig. 7, we report performance of our checkpoints, using datasets introduced in Fu et al. (2021) and commonly used in the literature. For each environment, we report the results on two different setting: (i) Low Budget, where the methods are given 25 0002500025\,00025 000 attempts, (ii) High Budget, where 100 000100000100\,000100 000 attempts are available. We also report results from concurrent RL methods (Qiu et al., 2022; Dernedde et al., 2024; Sun & Yang, 2023; Ye et al., 2023), without method-agnostic local search; and industrial solvers LKH3 (Helsgaun, 2017) and Concorde (Applegate et al., 2006).
First, we observe that stochastically sampling solutions with POMO for less than $10$ minutes already provides competitive results, ranking second among the five leading neural solvers. Adding EAS on top of POMO enables to compete with leading method MOCO on TSP500. This shows that auto-regressive RL constructive methods are already competitive at this scale. It is the first time that POMO is trained on those instance sizes, making it the first RL-trained neural solver that can be used for large instances both on TSP and CVRP, since most previous large scale RL methods are graph-specific, and hence cannot be applied on CVRP.
Furthermore, those results confirm that MEMENTO’s adaptation performance scales to instance size, and that it can still be zero-shot combined with COMPASS. On this benchmark, MEMENTO achieved state-of-the-art on $3$ of the $4$ tasks when zero-shot combined with COMPASS, showing significant improvement compared to previous state-of-the-art MOCO: the gap to optimality is reduced by more than $10\%$. It also proves efficiency on those large scales by being the best single agent method on all TSP instances, and on the low budget CVRP task.
@@ Results
Time performance at scale is key for the wide adoption of neural solvers since the number of sequential batches of attempts that can be achieved within a time budget will mostly depend on it (plus on the accessible hardware). Adaptation methods must be able to tackle large instances in reasonable time, and they must also scale well with the size of the base solver used, since Luo et al. (2023) demonstrated the importance of using large and multi-layered decoders for performance in neural CO (although they could not yet train them with RL).
Hence, we evaluate MEMENTO and EAS on a set of increasing instance sizes and decoder sizes and report their time complexity on the right part of Fig. 7. These plot show that, despite being slower for small settings, MEMENTO’s adaptation mechanism scales better than EAS. For an instance size of 1000, MEMENTO becomes $20\%$ faster to use the budget. And even for small instances (size $100$), using $10$ layers in the decoder’s architecture (1M parameters) makes EAS $40\%$ slower than MEMENTO. These scaling laws illustrate another benefit of using an approach that learns the update rather than relying on back-propagation to adapt neural solvers at inference time.

### v3

@@ 4 Experiments
We evaluate our method across widely recognized routing problems TSP and CVRP. These problems serve as standard benchmarks for evaluating RL-based CO methods (Deudon et al., 2018; Kool et al., 2019; Kwon et al., 2020; Grinsztajn et al., 2023; Hottung et al., 2022; Choo et al., 2022; Chalumeau et al., 2023b). In Section˜4.1, we validate our method by comparing it to leading approach for single-policy adaptation using online collected data, eas (Hottung et al., 2022). We present their comparative performance in line with established standard benchmarking from literature, and elucidate the mechanisms employed to adapt the action distribution of the base policy during inference. We then demonstrate how memento can be combined with SOTA neural solver compass with no additional retraining in Section˜4.2. Finally, we scale those methods to larger instances in Section˜4.3, outperforming heatmap-based methods.
Table 1: MEMENTO against baselines on a standard benchmark comprising 8 datasets with 4 distinct instance sizes on (a) TSP and (b) CVRP. Methods are evaluated on instances from the training distribution ($100$) and on larger instance sizes to test generalization. memento outperforms policy-gradient method eas and tree-search sgbs, with significant improvement over the base policy pomo. sgbs * results are reported from Choo et al. (2022) (bias discussed in Section˜A.1).
@@
We provide access to the code [^1] Code, checkpoints, and evaluation sets available at https://github.com/instadeepai/memento utilized for training our method and executing all baseline models. We release our checkpoints for all problem types and scales, accompanied by the necessary datasets to replicate our findings. We implement our method and experiments in JAX (Bradbury et al., 2018). The two problems are also JAX implementations from Jumanji (Bonnet et al., 2023). We use TPU v3-8 for our experiments.
@@ 4.1 Benchmarking MEMENTO against policy gradient fine-tuning
The most direct way to leverage online data for policy update is RL/policy-gradients. By contrast, memento learns an update based on the collected data. We therefore want to see how these two approaches compare. Since eas (Hottung et al., 2022) is SOTA method using policy gradient, we benchmark memento against it on a standard set of instances used in the literature (Kool et al., 2019; Kwon et al., 2020; Hottung et al., 2022). Specifically, for TSP and CVRP, these datasets comprise 10 00010\,000 instances drawn from the training distribution. These instances feature the positions of $100$ cities/customers uniformly sampled within the unit square. The benchmark also includes three datasets of distributions not encountered during training, each comprising $1000$ problem instances with larger sizes: $125$, $150$, and $200$, generated from a uniform distribution across the unit square. We employ the exact same datasets as those utilized in the literature.
@@
For routing problems such as TSP and CVRP, pomo (Kwon et al., 2020) is the base single-agent, one-shot architecture that underpins most RL construction solvers. In this set of experiments, memento and eas both use pomo as a base policy, and adapt it to get the best possible performance within a given budget. We train memento until it converges, on the same instance distribution as that used for the initial checkpoint. When assessing active-search performance, each method operates within a fixed budget of $1600$ attempts, a methodology akin to Hottung et al. (2022). In this setup, each attempt comprises of one trajectory per possible starting point. This standardized approach facilitates direct comparison to pomo and eas, which also utilize rollouts from all starting points at each step. We do not use augmentations with symmetries in our main results since they consume 87.5% of the budget to exploit network variance, which blurs the efficacy of the comparison and creates impractical settings for larger instances. Nevertheless, we report them in Section˜A.1 for continuity with previous literature.
@@ Results
The average performance of each method across all problem settings are presented in Table˜1. The observations we draw are three-fold. First, memento outperforms its base model (pomo) on the entire benchmark by a significant margin: showing that its adaptive search is superior to stochastic sampling. Second, memento outperforms eas on all $8$ tasks (spanning both in- and out-of-distribution) for both environments, highlighting the efficacy of learned policy updates when compared to vanilla policy gradients. Finally, we note that this improvement is significant; for example, on TSP100100, memento is doing 6060% better than sampling, while eas only 66%. We provide all time costs and performance comparison based on a time budget in Section˜A.1.
@@
To understand how memento uses its memory to derive a policy update, we analyse the logit update of an action with respect to its associated data, inspired by Lu et al. (2022); Jackson et al. (2024), and compare it to the REINFORCE policy gradient. On Fig.˜4, we plot the heatmap of the logit update with respect to the log probability (logp) and return associated with an action. We observe that the main rules learned by memento are similar to REINFORCE: an action with low logp and high return gets a positive update while an action with high logp and low return gets discouraged. It is interesting to see discrepancies. In particular, memento’s rules are not symmetric with respect to $x=0$, it only encourages action that are strictly above the mean return. A similar dissymetry would be expected with eas since it combines REINFORCE with a term that increases the likelihood to generate the best solution found. Appendix˜J extends this analysis by visualising the evolution of the update rule for different remaining budget.
Additionally, memento uses more inputs than REINFORCE: current budget available, trajectory logp, age of the data, and previous memento logit are also used to derive the new action logit. Appendix˜G provides an ablation study of these additional inputs, validating their importance.
Figure 4: Akin to REINFORCE (left), MEMENTO (right) encourages actions with high returns, particularly when they have low probability. memento learns an asymmetric rule: requiring the normalised return to be strictly positive to reinforce an action, but encouraging it even more.
4.2 Zero-shot combination with unseen solvers
In this section, we demonstrate zero-shot combination (no retraining) with compass. compass adapts by searching amongst a collection of diverse pre-trained policies.
Although providing fast adaptation, it can hardly improve further once the appropriate pre-trained policy has been found. In particular, online collected data cannot be used to update that policy’s action distribution. We use memento’s memory module, trained with pomo as a base policy, apply it on compass without additional retraining, and show empirically that we can switch on memento’s adaptation mechanism during the search and reach a new SOTA on $11$ out of $12$ tasks.
Figure 5: Combining MEMENTO and COMPASS during search on CVRP200, no re-training needed.
@@ 4.2 Zero-shot combination with unseen solvers
We combine our checkpoint of memento with a released model of compass (Chalumeau et al., 2023b). For the first half of the budget, we do not use memento, to let compass find the appropriate pre-trained policy. Once the search starts narrowing down, we activate memento to adapt the pre-trained policy, with no need to turn off compass’ search. Fig.˜5 shows the typical trends observed at inference. We recognise the typical search reported in Chalumeau et al. (2023b). At 5050% of budget consumption, memento is activated and one can observe a clear and significant drop in averaged tour length of the latest sampled solutions. The standard deviation also illustrates how the search is narrowed down.
Results
We evaluate the zero-shot combination of memento and compass on the standard benchmark reported in Section 4.1. Results for CVRP are presented in Table˜2, showing tour length and optimality gap in-distribution (CVRP100) and out-of-distribution. Although requiring no re-training, memento’s rules combine efficiently with compass and achieves new SOTA. Results on TSP are reported in Appendix˜A, and results on larger instances (TSP & CVRP) in Section˜4.3.
Table 2: SOTA performance on CVRP via zero-shot combination of memento and compass. The rules learned by memento transfer to the population-based method compass.
@@ 4.3 Scaling RL construction methods to larger instances
Scaling neural solvers to larger instances is a crucial challenge for the field. Auto-regressive construction-based solvers are promising approaches, requiring limited expert knowledge whilst providing strong performance. But are still hardly ever trained on larger scales with RL. As a result, their performance reported in the literature is usually quite unfair, being evaluated on large scales instances (>$500$) although having been trained on $100$ nodes (Qiu et al., 2022). In this section, we explain how we train construction solvers pomo and compass on instances of size $500$ with RL (on TSP and CVRP, checkpoints released); and use them as new base models to train and validate properties of memento at that scale.
@@
Training pomo on instances of size $500$ involves three main steps. First, inspired by curriculum-based methods, we start from a checkpoint pre-trained on TSP100. Second, we use Efficient Attention (Rabe and Staats, 2022) to reduce the memory cost of multi-head attention, enabling to rollout problems in parallel despite the $O(n^{2})$ memory requirement. Third, we use gradient accumulation to keep good estimates despite the constrained smaller instance batch size. These combined tricks enable to train pomo till convergence within $4$ days, and consequently train compass and memento. We also build the zero-shot combined memento (compass) checkpoint.
Figure 6: MEMENTO outperforms EAS on instances of size 500 across batch sizes and sequential attempts. Green areas indicate settings where memento adapts more efficiently. It consistently outperforms eas on TSP and in most CVRP settings, with strong gains under low budgets.
@@
On larger instances, the cost of constructing a solution is significantly increased (for pomo’s architecture: squared in computation and memory, and linear in sequential operations). It is also harder to get numerous parallel shots, which most methods depend on. It hence becomes crucial to develop data-efficient methods which stay robust to low budget regimes; and important to know which method to use for each budget constraint.
@@
We compare memento and eas on a set of 128 unseen instances (of size $500$), for a range of budget expressed in number of sequential attempts and number of parallel attempts attempts. Figure 6 reports the percentage of improvement (in absolute cost) brought by memento over eas. From this figure, we can observe three main tendencies. First, the data-efficiency of memento is illustrated by its consistent superiority on low budget (lower left of plots), reaching significant improvement compared to eas. Then, we can see that, on CVRP, the gap increases with the sequential budget for low parallelism. It demonstrates that memento’s update stays robust, whereas the gradient estimates of eas get deteriorated and cannot bring further improvement despite additional sequential budget. To finish with, we can see that memento fully dominates the heatmap on TSP, even for higher budget. However, when budget is increased for CVRP, eas is able to outperform memento, given large enough batches (>80>80).
Table 3: MEMENTO achieves SOTA performance on TSP500. It outperforms baselines in both budget regimes, attaining single-agent and overall SOTA when combined with pomo and compass, respectively.
Method Obj. Gap Time
Concorde 16.55 $0.000\%$ 38M
LKH-3 16.55 $0.003\%$ 46M
dimes 17.80 $7.55\%$ 2H
Difusco 17.23 $4.08\%$ 11M
DeepACO 18.74 $13.23\%$ -
moco 16.84 $1.72\%$ -
Pointerformer 17.14 $3.56\%$ 1M
Low Budget
@@
10M
High Budget
pomo (sampling)
eas
compass
memento (pomo)
memento (compass)
@@
29M
Results
In Table˜3 and Table˜4, we report performance of our checkpoints, using datasets introduced in Fu et al. (2021) and commonly used in the literature. For each environment, we report the results on two different setting: (i) Low Budget, where the methods are given 25 00025\,000 attempts, (ii) High Budget, where 100 000100\,000 attempts are available. We also report results from concurrent RL methods (Qiu et al., 2022; Dernedde et al., 2024; Sun and Yang, 2023; Ye et al., 2023), without method-agnostic local search; and industrial solvers LKH3 (Helsgaun, 2017) and Concorde (Applegate et al., 2006).
First, we observe that stochastically sampling solutions with pomo for less than $10$ minutes already provides competitive results, ranking second among the five leading neural solvers. Adding eas on top of pomo enables to compete with leading method moco on TSP500. This shows that auto-regressive RL constructive methods are competitive at this scale, contradicting previous literature (Qiu et al., 2022; Dernedde et al., 2024). It is the first time that pomo is trained at this scale, making it the first RL-trained neural solver that can be used for large instances on both TSP and CVRP: most previous large scale RL methods are graph-specific, and hence cannot be applied on CVRP.
These results confirm that memento scales to instance size, and can be zero-shot combined with compass. On this benchmark, memento achieved SOTA on $3$ of the $4$ tasks when zero-shot combined with compass, showing significant improvement compared to previous SOTA moco: the gap to optimality is reduced by more than $10\%$. It is the best single agent method on all TSP instances, and on the low budget CVRP task.
Table 4: MEMENTO outperforms baselines on CVRP500. It achieves single-agent SOTA on low budget regime with pomo and overall with compass.
Method Obj. Gap Time
LKH-3 37.229 $0.00\%$ 6H
Low Budget
@@ Results
12M
High Budget
pomo (sampling)
eas
compass
memento (pomo)
memento (compass)
@@
36M
@@
Time performance at scale is key to the adoption of neural solvers, because the number of sequential batches of attempts that can be achieved within a time budget will mostly depend on it. Search methods must be able to tackle large instances in reasonable time, and must scale well with the size of the base policy used, since Luo et al. (2023) demonstrated the importance of using large and multi-layered decoders in neural CO.
Hence, we evaluate memento and eas on a set of increasing instance sizes and decoder sizes and report their time complexity on Fig.˜1 (right). These plots show that, despite being slower for small settings, memento’s adaptation mechanism scales better than eas. For an instance size of 1000, memento becomes $20\%$ faster. And even for small instances (size $100$), using $10$ layers in the decoder’s architecture (1M parameters) makes eas $40\%$ slower than memento. These scaling laws illustrate another benefit of using an approach that learns the update rather than relying on back-propagation to adapt neural solvers at inference time.